# Foundation Models in Electrocardiogram: A Review

### v1

@@ 3.2 Pipeline
3.2.1 Pre-training Paradigm.
Pre-training represents a crucial first stage in developing ECG-FMs, as it equips the models with the knowledge necessary to adapt swiftly to various downstream tasks and contexts. The diverse modality of pre-training data, encompassing ECG signals, plots, and text-signal/plot pairs, allows for a wide spectrum of pre-training mechanisms in building and deploying ECG-FMs. In this survey, we introduce a new perspective that focuses on learning objectives during pre-training, categorizing existing methods for ECG-FMs into contrastive-based, reconstruction-based, generative-based, and hybrid techniques.
@@
Pre-training in ECG-FM studies generally involves the use of reconstruction techniques, such as masking and autoencoder-based methods. Masking, also known as masked representation learning, entails retaining a portion of the input and training the model to reconstruct the masked portion. Typically, this process begins by randomly masking a portion of the input ECG data, often by replacing it with a mask token or a zero-value, while keeping the rest of the data visible to the FM. For instance, adversarial masking [10, 11] involves using an image-based adversarial masking model to generate mask tokens as augmentations for each given ECG sample during pre-training, effectively reducing distortions of masked ECG signals. MassMIB [124], employs a random binary masking strategy on segmented patches of both raw ECG signals and the time-frequency spectrograms. BERTs and Masked Autoencoders (MAEs), along with their variants, are common masking-based pre-training methods in existing ECG-FM studies. BERT typically masks individual tokens, such as ECG timestamps or discrete channels, and leverages bidirectional context and self-attention to predict the masked portion through a Transformer encoder, while MAE masks larger chunks of the input, such as ECG plot patches or ECG series segments, and directly reconstructs the masked region using an encoder-decoder structure. By masking parts of the ECG data, the model can effectively learn to capture both local details (e.g., individual beats) and global structures (e.g., long-term heartbeat patterns), which is essential for robust representation learning.
@@
Autoencoder-based reconstruction is also essential for developing pre-trained ECG-FMs. Autoencoders are capable of learning a compressed representation of input ECG data by encoding it into a lower-dimensional latent space and then reconstructing the original ECG input from this compressed representation. Models like variational autoencoders (VAEs) and denoising autoencoders (DAEs) have demonstrated outstanding performance in numerous downstream tasks, including ECG synthesis and cardiac disease detection. As an example, Auto-TTE [16] harnesses a vector-quantized VAE to compress the raw ECG signals and quantize them into discrete latent representations, thereby enhancing the resilience of synthetic ECG data through the integration of compressed ECG information.
Generative-based pre-training.
@@
In generative-based pre-training, the FM generates future tokens by conditioning each token on the sequence of previously generated tokens. When applied to ECG analysis, this method typically utilizes Transformer-based models to encode input ECG data, such as heartbeat sequences and image patches, into high-dimensional tokens. These sequential tokens with position information enable FMs to capture temporal and spatial dependencies, as well as long-range relationships across ECG embeddings, like arrhythmias spanning multiple timestamps of recording. Current research commonly employs LLMs like GPTs as pre-trained encoders to obtain robust feature representations for downstream tasks. For instance, zero-shot RAG [127] uses LLaMA2 and GPT3.5 to convert ECG signals into prompts and integrates expert domain knowledge to guide LLMs in diagnosing arrhythmia and sleep apnea. The specific capability of generative-based pre-training methods on human instruction comprehension allows ECG-FMs to adapt to question-answering tasks, essentially acting as medical AI bots. SignalGPT [64], for example, offers an interactive interface for users to obtain essential information and fine-tunes ChatGPT by providing feedback. It uses ChatGPT as a comprehensive controller, determining the processing engine according to the signal types and analyzing the description of the given signal, based on pre-training on a massive ECG corpus.
@@
There are also several studies that utilize hybrid methods, combining contrastive-based, reconstruction-based, and generative-based techniques in the pre-training process. Hybrid pre-training has been shown to significantly improve the efficiency and robustness of representation learning by integrating the advantages of these methods. For example, GCL [99] integrates denoising autoencoder and augmentation contrast, where the decoder reconstructs the augmented views of ECG samples to reduce the noise introduced by data augmentation. CQA-ESI [128] adopts an LLM pipeline to generate detailed textual descriptions of ECGs and leverages contrastive and captioning loss to pre-train ECG encoders for enhanced representations.
3.2.2 Adaptation Paradigm.
@@
The ECG-FM adaptation paradigm customizes ECG-FM for specific tasks or datasets, thereby enhancing its performance through the utilization of acquired knowledge about cardiac data. Previous methods are classified into four branches: linear probe, fine-tuning, Low-Rank Adaptation (LoRA), and zero-shot inference.
@@
Linear probe is the simplest and most commonly used strategy for adapting ECG-FMs to downstream tasks. This approach involves training a linear classifier, such as logistic regression (LR), support vector machine (SVM), or random forest (RF), based on the feature representations learned by the pre-trained FM. During the adaptation process, only the weights of the linear classifier are updated through back-propagation, while the remaining FM parameters are kept frozen. The performance of this linear classifier can provide valuable insights into the usefulness of the learned representations for the task at hand. Consequently, it facilitates an assessment of the informational content within the model’s learned representations for a specific task or class, without the need for further fine-tuning of the model itself.
@@
Fine-tuning serves as a widely utilized strategy for adapting ECG-FMs, with two prevalent approaches: full fine-tuning and additive fine-tuning. Based on the way the foundation model is used on the target dataset, there are two mainstream works: full fine-tuning and additive fine-tuning. The former involves updating all parameters of the pre-trained model, while the latter selectively updates the parameters of newly introduced layers. Similar to linear probe, Most studies employ MLP or ResNet as the additive module. Full fine-tuning necessitates substantial task-specific data and computational resources, making it suitable for tasks significantly different from the initial training tasks, with ample task-specific data available. Conversely, additive fine-tuning offers a more practical option when computational resources are limited or the target task aligns closely with the pre-training tasks.
@@
[10] 11footnotemark: 1Adversarial masking ResNet1d-18 N/A Reconstruction Masking Linear LP CAC;
Age regression
PhysioNet/ CinC2020 Chapman ECG TS
[107] SleepFM CNN N/A Contrastive AC LR LP Sleep stage;
Sleep apnea
Multi-modal PSG Multi-modal PSG EEG;
ECG;
Respiratory
TS
[52] sEHR-ECG-Text BERT;
CNN;
GatorTron
EO Contrastive MC LR;
MLP
LP;
A-FT
CAC;
OOD
Mayo Clinic EHRs Mayo Clinic EHRs;
PhysioNet2020;
Chapman
ECG;
EHRs
TS+Text
[64] SignalGPT ChatGPT DO Generative AR MLP A-FT CAD;
Chatbot
CODE-test CODE-test ECG;
EEG;
EMG;
EOG
TS
[6] SimCLR;
BYOL;
VICReg
ResNet-50 N/A Contrastive AC ResNet1d-50 A-FT CAC CinC2020;
Ribeiro;
Chapman
PTB-XL ECG TS
[87] Contrastive SSL CNN1d N/A Contrastive AC MLP A-FT Stress WESAD;
RML
WESAD;
RML
ECG TS
[26] CardioGPT CLIP EO Contrastive MC Linear LP CAD GE&Philips;
NSH
GE&Philips;
NSH
ECG;
ECG reports
TF+Text
[53] DCL CNN1d
+Transformer
ED Contrastive AC Linear LP Emotion;
Blood glucose level
MAHNOB-HCI;
WESAD;
CASE;
PsPM-FR;
PsPM-RRM1-2;
ASCERTAIN
DREAMER;
D1NAMO
ECG TS
[61] ETP ResNet1d-18+ClinicalBERT EO Contrastive MC Linear LP;
ZS
CAC;
ZSC
PTB-XL;
CPSC2018
PTB-XL;
CPSC2018
ECG;
ECG reports
TS+Text
[29] Stacked Transformers Transformer ED Generative AR Linear;
LR
LP;
F-FT
CAC St-Petersburg;
PTB-XL;
Georgia;
Chapman;
Ningbo
CinC2021 ECG TS
[32] 3KG CNN1d N/A Contrastive AC Linear LP CAC PhysioNet2020 PhysioNet2020 ECG TS
[101] SimCLR;
BYOL;
SWaV
xResNet1d-50 N/A Contrastive AC Linear LP CAC PTB-XL;
Chapman
PTB-XL;
Chapman;
Ribeiro
ECG TS
[105] N/A EfficientNet-B3;
@@
N/A Contrastive AC Linear LP CAD PhysioNet/ CinC2021 PhysioNet/ CinC2021 ECG TS
[66] DLC AlexNet1d;
VGGNet1d;
ResNet1d
N/A Contrastive AC Linear LP;
F-FT
CAD NFH;
SNPH
PTB-XL;
CPSC2018;
Chapman
ECG TS
[74] CPC;
@@
[67] 22footnotemark: 2JCDCL CNN N/A Contrastive MC Linear LP CAD NFH PTB-XL;
CPSC2018
ECG TS+Plot
[41] ECG-MAE MLE-ViT EO Reconstruction Masking Linear LP;
F-FT
CAD CinC2020;
Chapman
PTB-XL ECG TS
[99] GCL ResNet1d;
VGGNet1d
N/A Hybrid AE+AC Linear LP;
F-FT
CAC;
Denoising
SNPH PTB-XL;
Chapman;
CPSC2018
@@
[96] MAESSL CNN;
ViT-Large
EO Reconstruction Masking MLP F-FT LVSD N/A N/A ECG TS
[46] ECG-GPT BEiT+GPT2 EO+DO Contrastive MC DistilBERT F-FT CAD YNHHS UK-Biobank;
PTB-XL;
Ribeiro
ECG;
ECG reports
Text+Plot
[83] 33footnotemark: 3Lead-agnostic SSL Wav2Vec+CMSC ED Contrastive AC Linear LP CAC;
@@
[124] 44footnotemark: 4MassMIB CNN+ViT EO Reconstruction Masking Linear F-FT CAD PTB-XL;
CPSC;
Chapman
PTB-XL;
CPSC;
Chapman
ECG TF
[51] 55footnotemark: 5N/A MSDNN+MoCo N/A Contrastive AC MSDNN F-FT CAD Cardiocloud W12L-ECG Cardiocloud W12L-ECG ECG TS
[110] 66footnotemark: 6HeartBEiT Dall-E+BEiT EO+DO Reconstruction Masking MLP F-FT LVEF;
HCM;
STEMI
MSHS MSHS;
PTB-XL
ECG Plot
[55] sCL-ST xResNet1d-50 N/A Contrastive AC Linear F-FT CAC CSPC;
INCART;
G12EC;
PTB
PTB-XL ECG TS
[16] 77footnotemark: 7Auto-TTE VQ-VAE DO Hybrid Reconstruction
+AR+GAN
N/A N/A Synthesis PTB-XL;
Sejong
PTB-XL;
Sejong
ECG;
ECG reports
TS+Text
[75] 88footnotemark: 8SSMs S4 with FCE N/A Contrastive PC MLP A-FT CAC CinC PTB-XL;
Chapman
ECG TS
[127] Zero-shot RAG LLaMA2;
GPT-3.5
DO Generative AR N/A LoRA;
ZS
CAC;
ZSC;
Sleep apnea
PTB-XL+;
Apnea-ECG
PTB-XL+;
Apnea-ECG
ECG;
ECG reports
TS+Text
[11] 99footnotemark: 9Adversarial masking ResNet1d-18 N/A Reconstruction Masking Linear LR CAC;
Gender
PhysioNet/ CinC2020 Chapman ECG TS
[135] MTECG Transformer ED Reconstruction Masking Linear A-FT AC Fuwai;
PTB-XL;
PCinC
Fuwai;
PTB-XL;
PCinC
ECG Plot
[1] LST EfficientNet1d;
ResNet1d;
ViT1d
EO Contrastive AC Linear LR Age;
BMI;
Sex
AHMS AHMS ECG;
PPG
TS
[80] ST-MEM ViT-B EO Reconstruction Masking Linear LP;
F-FT
CAC;
MIC
Chapman;
Ningbo;
CODE-15
PTB-XL;
CPSC2018;
PhysioNet2017
ECG Plot
[62] MERL-CKEPE ResNet1d-18+ Med-CPT DO Hybrid AR+Masking +MC Linear LP;
ZS
CAC;
MIC;
ZSC
MIMIC-ECG PTB-XL;
CPSC2018;
Chapman;
Ningbo
ECG;
ECG reports
TS+Text
[128] CQA-ESI GPT3.5;
ConvNext-v2;
BioLinkBERT
EO+DO Hybrid AR+MC Linear LP;
F-FT
CAC;
Patient;
Chatbot
PTB-XL;
Chapman;
MIMIC-IV-ECG
PTB-XL;
ICBEB
ECG;
ECG reports
TS+Text
[54] ISL ATT-CNN-RNN N/A Contrastive AC LR LP;
@@
[81] SimCLR CNN1d N/A Contrastive AC Linear LP Stress N/A N/A ECG TS
[113] ASTCL CNN1d
@@
ED Contrastive AC Linear LP;
@@
[68] MRC CNN N/A Contrastive AC Linear LP;
@@
[58] METS ResNet1d-18+ ClinicalBert EO Contrastive MC Linear LP;
ZS
ZSC MIMIC-III;
PTB-XL
PTB-XL;
MIT-BIH
ECG;
ECG reports
TS+Text
[130] MaeFE ViT1d;
Sub2D-ViT
EO Reconstruction Masking Linear LP;
F-FT
CAC CPSC2018;
PTB-XL;
Ningbo
PTB-XL;
CPSC2018
ECG TS+TF
[123] CNN+ViT Res2Net+SE
+ViT
EO Reconstruction Masking Linear F-FT CAC PTB-XL;
Chapman;
CPSC2018
PTB-XL;
Chapman;
CPSC2018
ECG TS
[85] SSL-KD SE-ResNet34 N/A Contrastive AC MLP F-FT CAC PhysioNet2020 PhysioNet2020 ECG TS+TF
[122] 1010footnotemark: 10BIOT Linear
@@
    AC Augmentation Contrast, MC Multimodal Contrast, AR Autoregressive-based,PC Prediction Contrast.
@@
    LP Linear Probe, F-FT Full Fine-tuning, A-FT Additive Fine-tuning, ZS Zero-shot.
@@
    CAC Cardiac Arrhythmia Classification, CAD Cardiac Abnormality Detection, MIC Myocardial Infarction Classification, Stress Stress Detection, Chatbot ECG Chatbot Assistant, Patient Patient Classification, Noise Noise Detection, ZSC Zero-shot Classification, Synthesis ECG Synthesis, Sleep stage Sleep Stage Classification, Sleep apnea Sleep Apnea Detection, OOD OOD Detection, Emotion Emotion Recognition, Blood glucose level Blood Glucose Level Monitoring, Denoising ECG Denoising, LVSD LVSD Diagnosis, LVEF LVEF Diagnosis, HCM HCM Diagnosis, STEMI STEMI Diagnosis, Gender Gender Classification, Age Age Classification, BMI BMI Classification, Sex Sex Classification.
@@
    NFH Ningbo First Hospital, SNPH Shanghai Ninth People’s Hospital.
@@
  • 2
    https://github.com/Aiwiscal/CDCL-ECG
  • 3
    https://github.com/Jwoo5/fairseq-signals
  • 4
@@
  • 5
    https://github.com/SMU-MedicalVision/ECG-Classfication
  • 6
    https://github.com/akhilvaid/HeartBEiT
  • 7
@@
  • 8
    https://github.com/tmehari/ssm_ecg
  • 9
    https://github.com/jessica-bo/advmask_ecg
  • 10
    https://github.com/ycq091044/BIOT

### v2

@@ 4.2 Pipeline
3.2.1 Pre-training Paradigms.
Pre-training represents a crucial first stage in developing ECG-FMs, as it equips the models with the knowledge necessary to adapt swiftly to various downstream tasks and contexts. The diverse modality of pre-training data, encompassing ECG signals, ECG plots, and text-signal/plot pairs, allows for a wide spectrum of pre-training mechanisms in building and deploying ECG-FMs. Notably, ECG plots here refer to visual views of ECG data, where raw signals are treated as 2d images. It is a trick widely used in ECG pre-processing, as it enables the application of advanced vision models to leverage the spatial-temporal structure of ECG data. In this survey, we introduce a new perspective that focuses on learning objectives during pre-training, categorizing existing methods for ECG-FMs into contrastive-based, reconstruction-based, predictive-based, and hybrid techniques.
@@
Pre-training in ECG-FM studies generally involves the use of reconstruction techniques, such as masking and autoencoder-based methods. Masking, also known as masked representation learning, entails retaining a portion of the input and training the model to reconstruct the masked portion. Typically, this process begins by randomly masking a portion of the input ECG data, often by replacing it with a mask token or a zero-value, while keeping the rest of the data visible to the FM. For instance, adversarial masking [10, 11] involves using an image-based adversarial masking model to generate mask tokens as augmentations for each given ECG sample during pre-training, effectively reducing distortions of masked ECG signals. MassMIB [132], employs a random binary masking strategy on segmented patches of both raw ECG signals and the time-frequency spectrograms. BERT along with their variants, are common masking-based pre-training methods in existing ECG-FM studies [118]. It typically masks individual tokens, such as ECG timestamps or discrete channels, and leverages bidirectional context and self-attention to predict the masked portion through a Transformer encoder.
@@
Autoencoder-based reconstruction is also essential for developing pre-trained ECG-FMs. Autoencoders are capable of learning a compressed representation of input ECG data by encoding it into a lower-dimensional latent space and then reconstructing the original ECG input from this compressed representation. Models like variational autoencoders (VAEs) and denoising autoencoders (DAEs) have demonstrated outstanding performance in numerous downstream tasks, including ECG synthesis and cardiac disease detection. As an example, Auto-TTE [16] harnesses a vector-quantized VAE to compress the raw ECG signals and quantize them into discrete latent representations, thereby enhancing the resilience of synthetic ECG data through the integration of compressed ECG information. Furthermore, masked Autoencoders (MAEs), represent a widely adopted approach in ECG reconstruction, combining a masking strategy with autoencoder architecture [131, 79, 44]. By masking portions of the input ECG signals or plots, MAEs reconstruct the masked regions using an encoder-decoder structure. This technique enables MAE to effectively capture both local details (e.g., individual beats) and global structures (e.g., long-term heartbeat patterns), which is essential for robust representation learning [103, 138].
Predictive-based pre-training.
@@
In predictive-based pre-training, the FM generates future tokens by conditioning each token on the sequence of previously generated tokens. When applied to ECG analysis, this method typically utilizes Transformer-based models to encode input ECG data, such as heartbeat sequences and image patches, into high-dimensional tokens. These sequential tokens with position information enable FMs to capture temporal and spatial dependencies, as well as long-range relationships across ECG embeddings, like arrhythmias spanning multiple timestamps of recording. Current research commonly employs LLMs like GPTs as pre-trained encoders to obtain robust feature representations for downstream tasks. For instance, zero-shot RAG [135] uses LLaMA2 and GPT3.5 to convert ECG signals into prompts and integrates expert domain knowledge to guide LLMs in diagnosing arrhythmia and sleep apnea. The specific capability of predictive-based pre-training methods on human instruction comprehension allows ECG-FMs to adapt to question-answering tasks, essentially acting as medical AI bots. SignalGPT [69], for example, offers an interactive interface for users to obtain essential information and fine-tunes ChatGPT by providing feedback. It uses ChatGPT as a comprehensive controller, determining the processing engine according to the signal types and analyzing the description of the given signal, based on pre-training on a massive ECG corpus.
@@
Several studies have effectively employed hybrid methods that combine contrastive-based, reconstruction-based, and predictive-based techniques in the pre-training process. Hybrid pre-training has been shown to significantly improve the efficiency and robustness of representation learning by integrating the advantages of these methods. First, hybrid pre-training can effectively mitigate the trade-off between sample diversity and the introduction of noise that generally accompanies data augmentation. For example, GCL [107] integrates denoising autoencoder and augmentation contrast, where the decoder reconstructs the augmented views of ECG samples to reduce the noise introduced by data augmentation. Furthermore, hybrid pre-training notably enhances the generalization capability of ECG-FMs by refining the multimodal representations of the samples. A compelling example is CQA-ESI [136], which adopts an LLM pipeline to generate detailed textual descriptions of ECGs and leverages contrastive and captioning loss to pre-train ECG encoders for enhanced representations. This approach demonstrates exceptional performance in zero-shot tasks.
3.2.2 Adaptation Techniques.
@@
The adaptation paradigm customizes ECG-FM for specific tasks or datasets, thereby enhancing its performance through the utilization of acquired knowledge about cardiac data. Previous methods are classified into four branches: linear probe, fine-tuning, Low-Rank Adaptation (LoRA), and zero-shot inference.
@@
Linear probe is the simplest and most commonly used strategy for adapting ECG-FMs to downstream tasks. This approach involves training a simple linear classifier, such as logistic regression (LR) or a fully-connected layer (FC), based on the feature representations learned by the pre-trained FM. During the adaptation process, only the weights of the linear classifier are updated through back-propagation, while the remaining FM parameters are kept frozen. The performance of this linear classifier can provide valuable insights into the usefulness of the learned representations for the task at hand. Consequently, it facilitates an assessment of the informational content within the model’s learned representations for a specific task or class, without the need for further fine-tuning of the model itself.
@@
Fine-tuning serves as a widely utilized strategy for adapting ECG-FMs, with three prevalent approaches: full fine-tuning, partial fine-tuning, and additive fine-tuning. Full fine-tuning involves updating all parameters of the pre-trained model, while partial fine-tuning first freezes the main part of the model and only finetunes the last layer or the last several layers. Additive fine-tuning only selectively updates the parameters of newly introduced layers. Similar to linear probe, most studies employ MLP or ResNet as the additive module. Full fine-tuning necessitates substantial task-specific data and computational resources, making it suitable for tasks significantly different from the initial training tasks, with ample task-specific data available [106]. Conversely, partial fine-tuning and additive fine-tuning offer a more practical option when computational resources are limited or the target task aligns closely with the pre-training tasks [93].
@@
Contrastive AC 3KG [34] CNN1d N/A FC LP CAC PhysioNet2020 PhysioNet2020 ECG TS
MLCAC [113] EfficientNet-B3;
@@
N/A FC LP CAD PhysioNet/ CinC2021 PhysioNet/ CinC2021 ECG TS
CPC;
@@
Contrastive SSL [95] CNN1d N/A MLP A-FT Stress WESAD;
RML
WESAD;
RML
@@
11footnotemark: 1Lead-agnostic SSL [90] wav2vec+CMSC ED FC LP CAC;
@@
ISL [59] ATT-CNN-RNN N/A LR LP;
@@
SSL-KD [92] SE-ResNet34 N/A MLP F-FT CAC PhysioNet2020 PhysioNet2020 ECG TS+TF
SimCLR;
BYOL;
VICReg [6]
ResNet-50 N/A ResNet1d-50 A-FT CAC CinC2020;
Ribeiro;
Chapman
PTB-XL ECG TS
SimCLR;
BYOL;
SWaV [109]
xResNet1d-50 N/A FC LP CAC PTB-XL;
Chapman
PTB-XL;
Chapman;
Ribeiro
ECG TS
DCL [57] CNN1d
@@
ED FC LP Emotion;
BGL
MAHNOB-HCI;
WESAD;
CASE;
PsPM-FR;
PsPM-RRM1-2;
ASCERTAIN
DREAMER;
D1NAMO
ECG TS
DLC [71] AlexNet1d;
VGGNet1d;
ResNet1d
N/A FC LP;
F-FT
CAD NFH;
SNPH
PTB-XL;
CPSC2018;
Chapman
ECG TS
22footnotemark: 2PIDA [55] MSDNN+MoCo N/A MSDNN F-FT CAD Cardiocloud W12L-ECG Cardiocloud W12L-ECG ECG TS
sCL-ST [60] xResNet1d-50 N/A FC F-FT CAC CSPC;
INCART;
G12EC;
PTB
PTB-XL ECG TS
LST [1] EfficientNet1d;
ResNet1d;
ViT1d
EO FC LR Age;
BMI;
Sex
AHMS AHMS ECG;
PPG
TS
SimCLR [88] CNN1d N/A FC LP Stress N/A N/A ECG TS
ASTCL [121] CNN1d
+Transformer
ED FC LP;
@@
Contrastive AC SleepFM [115] CNN N/A LR LP Sleep stage;
Sleep apnea
Multi-modal PSG Multi-modal PSG EEG;
ECG;
Respiratory
TS
MRC [73] CNN N/A FC LP;
@@
33footnotemark: 3BIOT [130] Linear
@@
44footnotemark: 4WCR [80] wav2vec+CMSC EO FC LP LVEF;
ACTN;
CAD
UHN-ECG;
PhysioNet2021;
MIMIC-IV-ECG
UHN-ECG ECG TS
ECGFounder [62] RegNet N/A FC LP CAC;
CAD;
CMD;
Demographics;
CED
HEEDB;
CODE-test;
PTB-XL;
PhysioNet2017
MIMIC-IV-ECG;
DeepBeat
ECG;
PPG
TS
MC sEHR-ECG-Text [56] BERT;
CNN;
GatorTron
EO LR;
MLP
LP;
A-FT
CAC;
OOD
Mayo Clinic EHRs Mayo Clinic EHRs;
PhysioNet2020;
Chapman
ECG;
EHRs
TS+Text
55footnotemark: 5JCDCL [72] CNN N/A FC LP CAD NFH PTB-XL;
CPSC2018
ECG TS+Plot
CardioGPT [27] CLIP EO FC LP CAD GE&Philips;
NSH
GE&Philips;
NSH
ECG;
ECG reports
TF+Text
ETP [66] ResNet1d-18+ClinicalBERT EO FC LP;
ZS
CAC;
ZSC
PTB-XL;
CPSC2018
PTB-XL;
CPSC2018
ECG;
ECG reports
TS+Text
ECG-GPT [50] BEiT+GPT2 ED DistilBERT P-FT CAD YNHHS UK-Biobank;
PTB-XL;
Ribeiro
ECG;
ECG reports
Text+Plot
METS [63] ResNet1d-18+ ClinicalBERT EO FC LP;
ZS
ZSC MIMIC-III;
PTB-XL
PTB-XL;
MIT-BIH
ECG;
ECG reports
TS+Text
CQA-ESI [136] ConvNext-v2;
BioLinkBERT
EO FC LP;
F-FT
CAC;
Patient;
Chatbot
PTB-XL;
Chapman;
MIMIC-IV-ECG
PTB-XL;
ICBEB
ECG;
ECG reports
TS+Text
PC 66footnotemark: 6SSMs [82] S4 with FCE N/A MLP A-FT CAC CinC PTB-XL;
Chapman
ECG TS
Reconstruction AE MAESSL [103] CNN;
ViT-Large
EO MLP F-FT LVSD N/A N/A ECG TS
MaeFE [138] ViT1d;
Sub2D-ViT
EO FC LP;
F-FT
CAC CPSC2018;
PTB-XL;
Ningbo
PTB-XL;
CPSC2018
ECG TS+TF
Reconstruction AE CNN+ViT [131] Res2Net+SE
+ViT
EO FC F-FT CAC PTB-XL;
Chapman;
CPSC2018
PTB-XL;
Chapman;
CPSC2018
ECG TS
ECG-MAE [44] MLE-ViT EO FC LP;
F-FT
CAD CinC2020;
Chapman
PTB-XL ECG TS
ECG-PCG-MAE [79] ViT EO FC LP SHDM;
AF;
LEF
AFib;
SHD;
Low-EF
AFib;
SHD;
Low-EF
ECG;
PCG
TS+FD
M 77footnotemark: 7Adversarial masking [10] ResNet1d-18 N/A FC LP CAC;
Age
PhysioNet/ CinC2020 Chapman ECG TS
77footnotemark: 7Adversarial masking [11] ResNet1d-18 N/A FC LR CAC;
Gender
PhysioNet/ CinC2020 Chapman ECG TS
88footnotemark: 8HeartBEiT [118] Dall-E+BEiT ED MLP F-FT LVEF;
HCM;
STEMI
MSHS MSHS;
PTB-XL
ECG Plot
MTECG [143] Transformer ED FC A-FT AC Fuwai;
PTB-XL;
PCinC
Fuwai;
PTB-XL;
PCinC
ECG Plot
99footnotemark: 9MassMIB [132] CNN+ViT EO FC F-FT CAD PTB-XL;
CPSC;
Chapman
PTB-XL;
CPSC;
Chapman
ECG TF
ST-MEM [87] ViT-B EO FC LP;
F-FT
CAC;
MIC
Chapman;
Ningbo;
CODE-15
PTB-XL;
CPSC2018;
PhysioNet2017
ECG Plot
Predictive AR Stacked Transformer [31] Transformer ED FC;
LR
LP;
F-FT
CAC St-Petersburg;
PTB-XL;
Georgia;
Chapman;
Ningbo
CinC2021 ECG TS
SignalGPT [69] ChatGPT DO MLP P-FT CAD;
Chatbot
CODE-test CODE-test ECG;
EEG;
EMG;
EOG
TS
Zero-shot RAG [135] LLaMA2;
GPT-3.5
DO N/A LoRA;
ZS
CAC;
ZSC;
Sleep apnea
PTB-XL+;
Apnea-ECG
PTB-XL+;
Apnea-ECG
ECG;
ECG reports
TS+Text
Hybrid AE+AR 1010footnotemark: 10Auto-TTE [16] VQ-VAE DO N/A N/A Synthesis PTB-XL;
Sejong
PTB-XL;
Sejong
ECG;
ECG reports
TS+Text
AE+AC GCL [107] ResNet1d;
VGGNet1d
N/A FC LP;
F-FT
CAC;
Denoising
SNPH PTB-XL;
Chapman;
CPSC2018
ECG TS
AR+M+MC MERL-CKEPE [67] ResNet1d-18+ Med-CPT DO FC LP;
ZS
CAC;
MIC;
ZSC
MIMIC-ECG PTB-XL;
CPSC2018;
Chapman;
Ningbo
ECG;
ECG reports
TS+Text
@@
    AC Augmentation Contrast, MC Multimodal Contrast, AC Augmentation Contrast, AR Autoregressive-based, M Masking, AE Autoencoder-based, PC Prediction Contrast.
@@
    FC Fully-connected Layer, LR Linear Regression, MLP Multilayer Perceptron.
@@
    LP Linear Probe, F-FT Full Fine-tuning, P-FT Partial Fine-tuning, A-FT Additive Fine-tuning, ZS Zero-shot.
@@
    CAC Cardiac Arrhythmia Classification, CAD Cardiac Abnormality Detection, MIC Myocardial Infarction Classification, Stress Stress Detection, BGL Blood Glucose Level Detection, Chatbot ECG Chatbot Assistant, Patient Patient Classification, Noise Noise Detection, ZSC Zero-shot Classification, Synthesis ECG Synthesis, Sleep stage Sleep Stage Classification, Sleep apnea Sleep Apnea Detection, OOD OOD Detection, Emotion Emotion Recognition, Blood glucose level Blood Glucose Level Monitoring, Denoising ECG Denoising, ACTN Abnormal Cardiac Troponin Diagnosis, LVSD LVSD Diagnosis, LVEF LVEF Diagnosis, HCM HCM Diagnosis, STEMI STEMI Diagnosis, SHDM Structural Heart Disease Murmurs Diagnosis, AF Atrial Fibrillation Diagnosis, LEF Low Ejection Fraction Diagnosis, Gender Gender Classification, Age Age Classification, BMI BMI Classification, Sex Sex Classification.
@@
    https://github.com/Jwoo5/fairseq-signals
  • 2
    https://github.com/SMU-MedicalVision/ECG-Classfication
  • 3
    https://github.com/ycq091044/BIOT
  • 4
    https://github.com/bowang-lab/ECG-FM
  • 5
    https://github.com/Aiwiscal/CDCL-ECG
  • 6
    https://github.com/tmehari/ssm_ecg
  • 7
@@
  • 8
    https://github.com/akhilvaid/HeartBEiT
  • 9
@@
  • 10